OpenStudy (anonymous):

Dallas solved the system of equations below. What mistake did he make in his work? 3x + y = 6 x − 2y = 2 y = 6 − 3x x − 2(6 − 3x) = 2 x − 12 + 6x = 2 5x − 12 = 2 +12 +12 x = 2.8 3(2.8) + y = 6 8.4 + y = 6 −8.4 −8.4 y = −2.4 He should have substituted 6 + 3x. He combined like terms incorrectly. It should have been 7x instead of 5x. He made a mistake in distributing −2. He made no mistake.

OpenStudy (jdoe0001):

so have you noticed anything wrong?

OpenStudy (jdoe0001):

$$ y = 6 − 3x\\ x − 2(6 − 3x) = 2\\ \color{red}{x} − 12 + \color{red}{6x} = 2\\ \color{red}{5x} − 12 = 2\\ +12 +12\\ \\ x = 2.8 $$ so, what do you think?

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw